# H-FA: A Hybrid Floating-Point and Logarithmic Approach to Hardware Accelerated FlashAttention

### v2

@@ III-B FlashAttention Units operating in parallel to multiple KV sub-blocks
Fig. 2 shows the parallel, block-level computation of the attention vector for a single query block. The query vector $\vec{q}$ is broadcast to multiple blocks, with each block handling a subset of the key-value matrix’ rows. Each FAU, for each sub-block, independently computes a partial result triplet $(\ell_{i},m_{i},\vec{o}_{i})$, which are then combined into the final output using cascaded ACC blocks that perform the merging described in Eq. (1). The FAU in each block is identical to the one in Fig. 1, except that the final normalization step (division) is deferred until all $KV$ blocks have been processed.
The operation of the parallel attention computation block is divided into two phases. In the first phase, all block-FAU units work in parallel to accumulate their respective $(\ell_{i},m_{i},\vec{o}_{i})$ triplets, each accessing all key–value vectors within a KV block. The time required for this phase is proportional to the depth of each KV block that equals to $N/p$, where $N$ refers to sequence length and $p$ to the number of parallel FAU units. In the second phase, the ACC units complete the accumulation of intermediate attention vectors through the vertical pipeline. To ensure synchronization between the two phases and smooth data transfer along the accumulation pipeline, the block-FAUs and ACC units communicate via a ready/valid pipelined flow-control protocol. Each ACC unit begins its computation once the outputs from both the block FAU and the preceding ACC are valid.
@@ IV-A Sum of two products in Logarithm Number System
The sign bit $s_{x}$ is 0 when $x$ is positive and $1$ otherwise. $X$ follows a fixed point representation with appropriately selected integer and fraction bits.
@@
According to [lns-add], based on the signs of $A$ and $B$, i.e., $s_{A}=s_{xz}$ and $s_{B}=s_{yu}$, we decide how to compute the LNS representation of $c$:
@@ V-A Logarithmic scaled output accumulation
First we remove in (14a) the need to compute logarithm by following Mitchell’s approximation [mitchell]. Since, by definition $0<2^{-|A-B|}<1$, Mitchell has shown in [mitchell] that $\log_{2}(1\pm 2^{-|A-B|})\approx 1\pm 2^{-|A-B|}$. Therefore, $\log_{2}|O_{i}|$ in (14a) can be equivalently written as:
@@ V-A2 Compute $B$
Up to now, we have shown how $A$ and $B$ are computed using only fixed point arithmetic. Thus, |AB||A-B| is also a fixed point result formed by the sum of its integer part $p$ and its fraction part $f$, i.e.,
@@ V-B Transforming fixed-point logarithm attention to a final floating-point attention result
To perform the needed transformation, first, we write the fixed point result $\text{log}_{2}|\text{Attn}(\vec{q},K,V)|$ as a sum of its integer $I$ and fractional $F$ part
@@
Since $F$ is a small number between 0 and $1$, we can utilize Mitchell’s approximation as in Eq. (18) but in the opposite direction to get that
@@
The right-hand side of (22) resembles the structure of a floating-point number, but without the bias in the exponent. To address this limitation, we add a bias term to $I$. Moreover, since $I$ can be either positive or negative, we adjust the sign of the final result accordingly, based on the outcome of (14d).
@@ VI-C Evaluation of hardware complexity
By increasing the number of parallel blocks we are able to achieve significant reduction in kernel execution time, by a factor of $6$ for $8$ parallel KV sub-blocks. Execution time considers only the latency needed to fetch the first query/ies, after that, queries are readily available, through pipelined memory accesses, when the accelerator is ready to start accumulating a new result. As we add more blocks the reaped performance benefits begin to plateau and will eventually come at a significant area overhead cost, since as shown in Fig. 8(b) area overhead keeps incrementing at approximately the same rate. For the case of $8$ parallel sub-blocks area is increased roughly $10\times$.

### v3

@@ III-B FlashAttention Units operating in parallel to multiple KV sub-blocks
Fig. 2 shows the parallel, block-level computation of the attention vector for a single query block. The query vector $\vec{q}$ is broadcast to multiple blocks, with each block handling a subset of the key-value matrix’ rows. Each FAU, for each sub-block, independently computes a partial result triplet $(\ell_{i},m_{i},\vec{o}_{i})$, which are then combined into the final output using cascaded ACC blocks that perform the merging described in Eq. (1). The FAU in each block is identical to the one in Fig. 1, except that the final normalization step (division) is deferred until all KVKV blocks have been processed.
The operation of the parallel attention computation block is divided into two phases. In the first phase, all block-FAU units work in parallel to accumulate their respective $(\ell_{i},m_{i},\vec{o}_{i})$ triplets, each accessing all key–value vectors within a KV block. The time required for this phase is proportional to the depth of each KV block that equals to $N/p$, where NN refers to sequence length and $p$ to the number of parallel FAU units. In the second phase, the ACC units complete the accumulation of intermediate attention vectors through the vertical pipeline. To ensure synchronization between the two phases and smooth data transfer along the accumulation pipeline, the block-FAUs and ACC units communicate via a ready/valid pipelined flow-control protocol. Each ACC unit begins its computation once the outputs from both the block FAU and the preceding ACC are valid.
@@ IV-A Sum of two products in Logarithm Number System
The sign bit $s_{x}$ is 0 when $x$ is positive and 11 otherwise. $X$ follows a fixed point representation with appropriately selected integer and fraction bits.
@@
According to [lns-add], based on the signs of AA and $B$, i.e., $s_{A}=s_{xz}$ and $s_{B}=s_{yu}$, we decide how to compute the LNS representation of $c$:
@@ V-A Logarithmic scaled output accumulation
First we remove in (14a) the need to compute logarithm by following Mitchell’s approximation [mitchell]. Since, by definition $0<2^{-|A-B|}<1$, Mitchell has shown in [mitchell] that $\log_{2}(1\pm 2^{-|A-B|})\approx\pm 2^{-|A-B|}$. Therefore, $\log_{2}|O_{i}|$ in (14a) can be equivalently written as:
@@ V-A2 Compute $B$
Up to now, we have shown how AA and $B$ are computed using only fixed point arithmetic. Thus, |AB||A-B| is also a fixed point result formed by the sum of its integer part $p$ and its fraction part $f$, i.e.,
@@ V-B Transforming fixed-point logarithm attention to a final floating-point attention result
To perform the needed transformation, first, we write the fixed point result $\text{log}_{2}|\text{Attn}(\vec{q},K,V)|$ as a sum of its integer II and fractional $F$ part
@@
Since $F$ is a small number between 0 and 11, we can utilize Mitchell’s approximation as in Eq. (18) but in the opposite direction to get that
@@
The right-hand side of (22) resembles the structure of a floating-point number, but without the bias in the exponent. To address this limitation, we add a bias term to II. Moreover, since II can be either positive or negative, we adjust the sign of the final result accordingly, based on the outcome of (14d).
@@ VI-C Evaluation of hardware complexity
By increasing the number of parallel blocks we are able to achieve significant reduction in kernel execution time, by a factor of 66 for 88 parallel KV sub-blocks. Execution time considers only the latency needed to fetch the first query/ies, after that, queries are readily available, through pipelined memory accesses, when the accelerator is ready to start accumulating a new result. As we add more blocks the reaped performance benefits begin to plateau and will eventually come at a significant area overhead cost, since as shown in Fig. 8(b) area overhead keeps incrementing at approximately the same rate. For the case of 88 parallel sub-blocks area is increased roughly $10\times$.